OpenStudy (anonymous):

convert 7π÷13 to degree

OpenStudy (anonymous):

\[2\Pi= 360^{0}\] \[(7\Pi \div 13)\times (360^{0}\div2\Pi)= 96.9230769 degree\]

OpenStudy (wolf1728):

7 PI / 13 = 1.6916268135 radians 7 PI / 13 * (180/PI) = 7 / 13 * (180) = 96.9230769230769 degrees